### v1

# A universal neutral-atom quantum computer with individual optical addressing and non-destructive readout

Abstract
Quantum computers must achieve large-scale, fault-tolerant operation to deliver on their promise of transformational processing power [1, 2, 3, 4]. This will require thousands or millions of high-fidelity quantum gates and similar numbers of qubits [5]. Demonstrations using neutral-atom qubits trapped and manipulated by lasers have shown that this modality can provide high two-qubit gate (CZ) fidelities and scalable operation [6, 7, 8, 9, 10]. However, the gates in these demonstrations are driven by lasers that do not resolve individual qubits, with universal computation enabled by physical mid-circuit shuttling of the qubits. This relatively slow operation will greatly extend runtimes for useful, large-scale computation. Here we demonstrate a universal neutral-atom quantum computer with gate rates limited by optical switching times, rather than shuttling, by individually addressing tightly focused laser beams at an array of single atoms. We achieve CZ fidelity of 99.35(4)and local single-qubit $R_{Z}$ gate fidelity of 99.902(8)%. Moreover, we demonstrate non-destructive readout of alkali-atom qubits with sub-percent loss, which boosts operational speed. This technique also enables us to measure 99.73(3)CZ fidelity with atom-loss events excluded, which is a record among long-lived neutral-atom qubits and highlights the path to higher fidelity and error correction. Our results represent a critical step towards large-scale, fault-tolerant neutral-atom quantum computers that can execute computations on practical timescales.
1 Introduction
Quantum computers promise to reduce computational energy consumption and solve classically expensive or intractable problems in a range of fields including material science, machine learning, security, chemistry, energy, finance, medicine, and pharmaceutical development [1, 2, 3, 4, 11, 12, 13, 14, 15]. Recent analyses suggest that realization of practical utility will require fault-tolerant operation [5]. This in turn requires quantum error correction (QEC), which comes with large overhead in the number of qubits and requires high one- and two-qubit quantum gate fidelities. The field’s progress along this path has entered a phase of experimental demonstrations [16, 17, 7]. Among the diverse modalities proposed for large-scale fault-tolerant quantum computation [18], the neutral-atom approach is one of the most promising candidates and has become a leader in the combination of qubit count, universal gate-based circuit operation, connectivity, and quantum error detection [7, 19, 20]. The rapid progress of neutral atoms for quantum computing can be attributed to the combination of naturally identical qubits; long coherence times due to decoupling from the environment [21]; switchable qubit interaction strengths across 12 orders of magnitude [22]; sub-microsecond, high-fidelity, scalable gates [6, 23]; and the ability to trap and manipulate large, dense, multi-dimensional qubit arrays [24, 19, 25], as was recently highlighted with a demonstration of error-detected logical qubits [7].
Two architectural approaches have emerged for neutral-atom quantum processors. One [7] relies on simultaneously shuttling multiple qubits to and from a single, dedicated zone for implementing gates. This allows flexible qubit connectivity and avoids the need for near-diffraction-limited addressing beams but increases compilation complexity [26] and limits gate rates due to constraints on the speed of atom movement[27]. High gate rates are important for maintaining practical computation times, particularly in light of the physical qubit count and gate depth overheads required for fault-tolerant operation. An alternative approach [20] keeps the qubits stationary and addresses them individually with steerable laser beams, enabling gate rates limited only by optical switching, which is inherently faster than atom movement. This stationary qubit approach poses challenges for achieving high gate fidelities and does not natively support all-to-all connectivity, but it offers the promise of faster circuits, in direct support of long-term fault-tolerant operations [28]. Here, we present a full-stack neutral-atom quantum computer using gates based on individual optical addressing instead of atom motion. Overcoming the challenges specific to this architecture, we achieve entangling gate fidelity of 99.35(4)%, already exceeding QEC thresholds for certain codes [29, 30, 31, 32]. Our results open the door to faster approaches for fault-tolerant operation that exclusively use individual optical addressing of stationary qubits, or that combine this with some atom shuttling to optimize execution of error-corrected circuits. We further advance the state-of-the-art by incorporating non-destructive qubit readout, immediately enabling faster computations and establishing a key ingredient for future mid-circuit measurements and error correction.
2 A neutral-atom architecture with individual optical addressing
Figure 1: System overview. a, Image of array of Cs atoms, trapped in a glass cell. b, Locally addressed $R_{Z}(\phi)$ (beam shown in blue) and CZ (beams show in purple) gates are applied via beams propagating perpendicular to the atom plane. Microwave-based global $R(\theta,\phi)$ gates (labelled GR) address the entire array. c, Layout of key lasers implementing trapping and gates. Rydberg excitation is implemented by counter-propagating 459 nm (blue) and 1040 nm (red) lasers. Rydberg illumination for control and target qubits are sourced from separate pulse shaping systems (labeled A and B). An optical tweezer array (1064 nm) is generated via AODs and combined with the 1040 nm light. Readout fluorescence is collected and imaged on an electron-multiplying CCD (EMCCD). d, Atomic levels and addressing fields, including lasers for two-photon Rydberg excitation, readout, and optical pumping (OP). Rydberg beams are circularly polarized ($\sigma^{+}$ and $\sigma^{-}$ for 459 nm and 1040 nm, respectively). Rydberg blockade shift denoted by $V$.
Building on prior work [20], we present a full-stack neutral-atom quantum computer using an individual optical addressing architecture. Arbitrary quantum circuits can be submitted via the cloud, where they are compiled333via Infleqtion Superstaq [33] into a universal native gate set consisting of global microwave $R(\theta,\phi)$ rotations, arbitrary local $R_{Z}(\phi)$ rotations, and nearest-neighbor controlled-Z (CZ) gates. The qubits are optically trapped Cs atoms, with |$\ket{0}$ and |$\ket{1}$ encoded in the |$\ket{F=3,m_{F}=0}$ and |$\ket{F=4,m_{F}=0}$ clock states of the ground state hyperfine manifold, respectively. The atoms are trapped out of a cold atomic beam [^4] Infleqtion PICAS, cooled, and rearranged by an optical tweezer to form a defect-free array. Fig. 1 illustrates a 24-qubit computational array alongside 12 reservoir atoms, formed in a 1064 nm tweezer array (see Methods Sec. 6.1). After state preparation via optical pumping, arbitrary circuits are implemented with microwave and laser pulses, followed by non-destructive state-selective readout (NDSSR) of the qubit state. During this readout the qubits remain trapped in the optical tweezers with high probability, allowing re-initialization and re-use of the same atomic array for the next shot (see Sec. 4). After the requested number of measurements is collected, circuit results are returned to the user. Throughout this process feed-forward stabilization is applied to background electric and magnetic fields, and lasers are stabilized on various timescales using both optical power measurements and atom-based signals.
Unique to this system is the optical design supporting laser-based local $R_{Z}(\phi)$ rotations and CZ gates. The former requires application of 459 nm light to a single atom, and the latter requires simultaneous application of 459 nm and 1040 nm light to a pair of atoms to drive two-photon Rydberg excitation. To implement these gates, we send each laser through a sequence of three acousto-optic devices: one for pulse shaping and two for beam steering. The pulsed and steered beams are then focused to nearly diffraction-limited spots on the atom array (see Fig. 1). Using both direct optical measurements and atom-based beam profiling, we confirm that our 459 and 1040 beams are focused to spot sizes of 2.8 $\mu\rm m$ and 3.0 $\mu\rm m$ ($1/e^{2}$ intensity radii) in the atom plane. Critically, these spots are smaller than the atom spacing of 6 $\mu\rm m$, allowing us to address all qubits individually. With this approach, we can alternate between CZ and local $R_{Z}(\phi)$ gates (on select target qubits) with a timescale limited only by the beam-steering devices. Our current acousto-optic deflectors (AODs) used as beam-steering devices can be re-aimed with a timescale of 2 $\mu\rm s$. In future devices, combining AODs with spatial light modulators (SLMs) can enable fast, individually addressed gates on multiple qubits simultaneously [34]. With this approach our individual-addressing architecture supports parallel gate operations while maintaining high gate rates.
A practical consideration for this individual addressing design is the crosstalk that can occur from residual beam intensity reaching a neighboring atom. For instance, given our beam waists, one predicts 0.03% intensity reaching neighboring qubits. In practice, we find crosstalk intensities up to ${\sim}10$x higher, likely due to optical aberrations. We provide a thorough discussion of the implications of this crosstalk in Methods Sec. 6.7, and summarize the results here. One effect of this crosstalk is residual light reaching spectator atoms not participating in the gate. For our current parameters, we estimate that this produces an error contribution smaller than $10^{-4}$ per gate. Moreover, it is a predictable unitary error that can be compiled away [33]. The other important impact of crosstalk is on the entangling gate itself, where interference between the control and target illumination can significantly perturb the gate dynamics. However, we find that this error contribution can be made negligible by ensuring that the fields interfere at a frequency much higher than the Rabi frequency of the Rydberg excitation. We design our addressing system to satisfy this requirement. For each wavelength, we branch the single source laser into two fully independent beam-shaping and beam-steering paths (labeled A and B in Fig. 1). Thus, in a CZ gate, the Rydberg laser light reaching the control and target qubits have disparate acousto-optic frequency shifts, resulting in controllable crosstalk beatnote frequencies in the range $0-30$ MHz (see Methods Sec. 6.7). This dual beamline system enables our processor to tolerate greater crosstalk intensity at higher array density, achieving higher blockade strength, while still addressing qubits individually and minimizing the impact of crosstalk on neighboring atoms.
3 Quantum processor implementation and characterization
Optically trapped Cesium atoms offer a hyperfine qubit with excellent coherence properties. We find a trap vacuum lifetime of 26(2) s, limited by background gas collisions, and measure an atom temperature of 2.6(3) $\mu$K through a combination of release-recapture[35] and ground-Rydberg coherence measurements. This temperature alone would permit a qubit coherence time of about 42 ms. In practice we measure $T_{2}^{*}=12.7(5)$ ms, and we attribute this additional decoherence to magnetic field shot-to-shot fluctuation of ${\sim}1$ mG at our bias of 11.15 G, as supported by measurements of $T_{2}^{*}<1\,\text{ms}$ on magnetically sensitive hyperfine transitions. By concatenating repeated XY8 dynamical decoupling sequences, we can extend $T_{2}$ beyond the capability of our hardware to accurately measure it—our best estimate is 2.8 s, with a 97% confidence interval lower bound of 128 ms.
Our single-qubit gate set consists of global $R(\theta,\phi)$ (GR) and local $R_{Z}(\phi)$ rotations. The former are implemented via a uniform pulsed microwave field resonant with the magnetically insensitive qubit |$\ket{0}\rightarrow\ket{1}$ transition. The duration and phase of the pulse determines $\theta$ and $\phi$, respectively. Fig. 2a illustrates Rabi oscillations driven by this microwave field. With this system, we have reached Rabi rates as high as $\Omega_{\mathrm{GR}}$ $=2\pi\times 120$ kHz, enabling an $R(\pi,\phi)$ gate time of 4.1 $\mu$s. We characterize the full set of GR Cliffords across a 24-qubit system using randomized benchmarking (Fig. 2b, d), and find an average fidelity of 99.980%, with an ensemble standard deviation of 0.003% (Fig. 2) and a typical fit uncertainty of 0.004%.
Local $R_{Z}(\phi)$ rotations are implemented via the differential light shift arising from a locally addressed, 459 nm laser beam that off-resonantly addresses the transition from |$\ket{1}$ to the $7P_{1/2}$ excited state with a detuning $\Delta$ from the $7P_{1/2}$ center of mass. Fig. 2b presents Ramsey measurements of a differential light shift of $2\pi\times 2$ MHz, enabling an $R_{Z}($$\pi$)))) gate time of 250 ns. We characterize $R_{Z}(\phi)$ fidelity for two values of $\Delta$ using a novel form of interleaved randomized benchmarking (IRB), which samples rotation angles uniformly distributed over $[0,2\pi)$ (see Methods Sec. 6.3). We first characterize the full 24-qubit array with $\Delta/2\pi=885$ MHz. This characterization uses a maximum IRB depth of $200$ $R_{Z}(\phi)$ gates per qubit, which entails redirecting the beam-steering AODs $4800$ times in a 15 ms circuit. We find an average $R_{Z}(\phi)$ fidelity of 99.76%, with an ensemble standard deviation of 0.03% (Fig. 2e). To properly account for the difference between leakage and true depolarization, the fidelity calculation includes an estimate of the scattering error due to non-zero population of the $7P_{1/2}$ intermediate state (Methods Sec. 6.3). Our observed $R_{Z}(\phi)$ error is dominated by this scattering, which can be suppressed by using larger detuning. Indeed, changing $\Delta/2\pi$ to $-2.1$ GHz, we measure, for a single qubit, a higher $R_{Z}(\phi)$ fidelity of 99.902(8)(see Fig. 2d). The residual error of 0.098(1)is largely explained by the remaining leakage probability of 0.066% and laser pulse energy fluctuation.
Figure 2: Single-qubit gates. a, Rabi oscillations driven by a global microwave field. b, Ramsey oscillations from local qubit $R_{Z}(\phi)$ control. c, d, Randomized benchmarking results for global $R(\theta,\phi)$ and local $R_{Z}(\phi)$ gates, respectively. e, Randomized benchmarking results, plotted as cumulative distribution function (CDF) of 24-qubit (or 1-qubit) results. For 24-qubit results, gates are applied to each of the 24 qubits in each circuit.
Our quantum processor implements the two-qubit controlled-Z (CZ) gate, which, combined with arbitrary single-qubit rotations, is sufficient for universal quantum computation. The gate is based on simultaneous excitation of the control and target qubits to a Rydberg state, where blockade interactions provide inter-qubit coupling. In particular, we use two-photon excitation (459 nm + 1040 nm) to couple the |$\ket{1}$ state to the |$\ket{69S_{1/2},m_{j}=-1/2}$ Rydberg level. The single-photon intensities are defined by the $\Delta/2\pi=-2.1$ GHz detuning from the $7P_{1/2}$ intermediate state, the $2\pi\times 2$ MHz light shift applied to the qubit by the 459 nm laser and the $\Omega=2\pi\times 3$ MHz Rabi frequency of Rydberg excitation (Fig. 3a). We measure a blockade interaction energy of $2\pi\hbar\times 10.3(1)$ MHz (Fig. 3b). We implement an approximation of the time-optimal Rydberg CZ gate [23] using a pulse of duration 7.85/$\Omega$ = 416 ns with a sinusoidal phase profile (implemented via electro-optic modulation of the 1040 nm beam; see Fig. 3c). The full CZ gate consists of this phase-modulated Rydberg pulse followed by single-qubit phase corrections to produce the canonical gate. We first calibrate the uncorrected gate, using a cost function that is insensitive to the single-qubit phases [6] (Fig. 3d), then calibrate the phase corrections (see Methods Sec. 6.4).
We characterize our CZ gates using randomized benchmarking circuits in which CZ gates alternate with global single-qubit gates (see Fig. 3e). Symmetric benchmarking is discussed in detail in Ref. [36], and approaches similar to ours were used in Refs. [6, 37, 8]. This approach has limitations; for example, swap errors go completely undetected, and it strictly only characterizes the performance of the gate on a subset of two-qubit input states. Nevertheless, it is known to perform well given the error profile of neutral-atom platforms. To validate application of the approach to our gate, we perform density matrix simulations of our benchmarking procedure using a realistic microscopic error model for the gate (see Methods Sec. 6.6) and find agreement between the fidelity extracted from benchmarking and the true gate fidelity averaged over Haar-random input states, reproducing findings from prior applications of the technique [6, 37].
Figure 3 presents CZ benchmarking results for a single pair of atoms. In these circuits single-qubit phase corrections are implemented via a combination of local $R_{Z}(\phi)$ and virtual phase shifts to the local oscillator, as in Ref. [38]. Because we characterize a single qubit pair, the beam-steering AODs are statically driven throughout the control sequence. By following each NDSSR measurement immediately with an occupancy measurement, we are able to separately characterize the two-atom survival probability ($R=99.62(3)$per gate) and cycle polarization ($P=99.67(4)$%). Combining these with a model-based estimate of the leakage probability (see Methods Sec. 6.5), we extract a CZ fidelity of $99.35(4)$%. The leading physical mechanisms contributing to the observed $0.65$% CZ gate error are leakage due to the finite lifetime of the $7P_{1/2}$ intermediate state, leakage and loss due to the finite lifetime of the Rydberg level, and shot-to-shot intensity fluctuations of the Rydberg laser beams. We estimate the contribution from optical crosstalk to CZ error to be $\lesssim 0.1$% in our current configuration, with a straightforward path to further reduction (see Methods Sec. 6.7).
NDSSR allows us to filter the data based on atom retention and calculate a post-selected CZ fidelity of 99.73(3)(see Methods Sec. 6.5). This ${\sim}0.38$% atom loss error may be converted to erasure through a leakage-detection measurement in which an ancilla becomes entangled with the presence/absence of an atom in a qubit state (rather than its logical state) [39], enabling the use of erasure-specific techniques to achieve improved correction performance [40].
Figure 3: Characterization of CZ gates. a, Rabi oscillations between |$\ket{1}$ and |$\ket{R}$, at a rate of $2\pi\times 3$ MHz. b, Rydberg blockade spectroscopy. c, Measured phase profile of 1040 nm optical pulse used in the CZ gate. d, Circuit used as cost function for CZ gate optimization. Global $R(\theta,\phi=0)$ gates denoted as $X_{\theta}$. e, Circuits used for characterization of CZ gates include a main block (alternating Haar-random global SU(2) gates and CZ gates), as well as an inversion operation (consisting of three CZ gates and four global rotation gates $R_{i}$) and a padding circuit that reduces the spurious impact of $R(\theta,\phi)$ errors on extracted CZ fidelity. For circuits with $N=0$ the “Main” and “Inversion” blocks are omitted entirely. f, Observed probability to retain both atoms as a function of the number of CZ gates in the circuit, along with the fit used to extract the retention probability. Decay is fit to an asymptote of zero. g, Observed probability to obtain the expected state-selective measurement outcome, given that both atoms were retained, as a function of the number of CZ gates in the circuit, along with the extracted cycle polarization. Decay is fit to an asymptote of 25%. Data in f, g are obtained from sequential NDSSR (g) and occupancy readout (f) measurements on the same quantum state. Each point is obtained as an average over between 18 and 48 circuits (183 total circuits in dataset), with an error bar corresponding to standard error of the mean.
4 Non-destructive State Selective Readout
While neutral-atom platforms have demonstrated significant improvements in gate fidelity and qubit count scaling that make neutral atoms a leading modality for quantum computing, there has been relatively less emphasis on their shot rate. Non-circuit operations such as atom loading and arrangement are expected to be a non-negligible part of the total operation time for measurement schemes based on destructive readout; a non-destructive readout scheme with repetitive measurements can significantly increase shot rate by amortizing the cost of these non-circuit operations and can enable mid-circuit measurements for fault-tolerant operation.
We implement the non-destructive state detection method demonstrated in Ref. [41] with a few critical improvements (see Methods 6.8). During the state detection sequence, qubit state $|1\rangle$ is optically pumped to the stretched state |$\ket{F=4,m_{F}=4}$, where we can detect fluorescence from the closed |$\ket{F=4,m_{F}=4}\rightarrow\ket{F^{\prime}=5,m_{F}^{\prime}=5}$ transition, while |$\ket{0}$ remains dark. By strategically alternating between illumination with trap light and readout light, we mitigate the deleterious effects of excited state light shift and dipole force fluctuations, and we reach a state-averaged loss rate of $0.86(25)\%$ while achieving bright-dark discrimination fidelity of 99.6(2)%with 6 ms of camera exposure time. The imaging performance and the state-discrimination fidelity of the readout scheme is analyzed from a large dataset of images collected during state preparation and measurement (SPAM) characterization experiments, each of which prepares all qubits to be in $|1\rangle$ (bright) or $|0\rangle$ (dark) (Fig. 4). The raw state-discrimination fidelity without state preparation or atom loss correction is 97.7(5)%, limited by bright state detection error of 3.1(7)and dark state detection error of 1.6(5)%. Note that state preparation error, which includes failure to optically pump into $|F=4,m_{F}=4\rangle$ during Raman sideband cooling and failure to coherently transfer $|F=4,m_{F}=4\rangle$ to $|F=4,m_{F}=0\rangle$ using microwave pulses (see Sec. 6.2), is distributed across both types of state detection errors.
The shot rate enhancement from repetitive non-destructive measurements is demonstrated by recording 8.2(2) Hz shot rate for a 15-qubit Greenberger–Horne–Zeilinger (GHZ) state preparation circuit executed within a 24-qubit array (see Sec. 6.8.3), compared to an estimated single-measurement shot rate of ${\sim}3$ Hz. These experiments were performed without Raman sideband cooling, the inclusion of which adds additional time to each measurement that cannot be amortized (see Methods Fig. 9). To the best of our knowledge, the only other demonstration of repeated non-destructive measurements of neutral-atom-based quantum circuits utilizing a universal gate set is found in Ref. [9]. This demonstration used an alkaline-earth species (strontium-88) with 10 ms of exposure time for non-destructive measurement of ancilla atoms with 3.5% atom loss rate.
The non-destructive readout method for alkali atomic species demonstrated in this work provides the starting point for repetitive mid-circuit readout of ancilla qubits in dual-alkali-species-based logical qubit architectures. The atom loss rate can be further reduced by using shorter exposure times and advanced camera image binarization algorithms that work at low photon counts [42]. The depumping error during imaging can be improved with technical improvements in polarization purity and with microwave-assisted state-selective repumping during imaging [43].
Figure 4: Non-destructive state-selective readout. NDSSR signal distribution from 51862 images collected during 43 iterations of SPAM characterization. The turquoise-colored distribution describes the signals where atoms are expected to be in a bright state ($|F=4\rangle$) and the purple-colored distribution describes the signals for a dark state ($|F=3\rangle$). The samples for each distribution come from the deterministically loaded diagnostic site pair used for CZ gate fidelity benchmarking. Photo-electron counts for each site were calculated from the raw EMCCD camera analog-digital units using nominal EM gain and pre-amplifier gain. Each quadrant of the plane defined by NDSSR and post-NDSSR photo-electron counts contains two text labels showing total counts from $|F=4\rangle$ preparation and measurement experiment (turquoise) and counts from $|F=3\rangle$ preparation and measurement experiment (purple). a, Sampled probability distribution of photo-electron counts for bright and dark states. The dashed black line denotes the discrimination threshold that yields discrimination fidelity of 99.6(2)%. The shaded areas under the distributions on both sides of the discrimination threshold denote signals associated with state detection error. Depumping and atom loss during the NDSSR process and microwave pulse infidelity during state transfer from the stretched state to the bright qubit state give rise to bright state detection error of 3.1(7)%, which is reduced to 2.6(7)when atom loss is post-selected out. The dark state detection error of 1.6(5)is believed to be a combination of (partial) state preparation error and $F=3\rightarrow F=4$ repumping during imaging from residual leakage of repump laser intensity. The middle inset shows a single NDSSR measurement of a four-qubit GHZ state realized in the 24-qubit array. b, NDSSR and post-NDSSR occupancy readout photo-electron counts. The bright state atom loss rate is 1.0(4)and the dark state atom loss rate is 0.8(3)%. The atom loss rates and state detection errors provided are the means and standard deviation of their measured values for the 43 experimental iterations that comprise the dataset.
5 Outlook
This work advances the state-of-the-art for individually addressed entangling gates in a neutral-atom platform, and we foresee a clear path towards the error levels required for fault tolerance. Control errors may be reduced by reaching lower atom temperatures, better stabilizing laser power and frequency, and using robust gate protocols[44, 45]. Straightforward changes such as increasing the intermediate-state detuning will immediately reduce scattering error. Moving forward, scattering error can be eliminated by transitioning to single-photon Rydberg excitation [38, 37], which also reduces the complexity of light shifts on atomic resonances. The path for scaling our qubit array is also well-established, with demonstrations[46] suggesting the feasibility of trapping $>10,000$ qubits in a single array, with the option to improve trapping lifetimes by operating at cryogenic temperatures[47]. The non-destructive readout we present here already enables state-of-the-art shot rate, and paths towards even faster readout are being developed[48, 49]. Finally, we note that our ability to achieve high ground-Rydberg coherence ($T_{2}^{*}=15~{}\mu\text{s}$) at $n=69$ is a unique advantage in the field, enabled by our electric field control. Since the Rydberg interaction scales with $n^{11}$, operating at higher $n$ unlocks greater connectivity, which is valuable for QEC protocols and applications[28].
For quantum computers to provide meaningful scientific and commercial value, they will likely need to support fault-tolerant operation on at least $\sim$100 logical qubits[50, 51, 5], and neutral-atom architectures are evolving rapidly towards that goal. This work demonstrates important new capabilities on that path, namely the ability to implement individually addressed high fidelity entangling gates and non-destructive state-selective readout, which collectively enable faster qubit operations and a broader set of QEC protocols. Ultimately, a compelling architecture for fast, fault-tolerant quantum computing is likely to integrate these techniques alongside other demonstrated neutral-atom capabilities, including dual-species platforms[52, 53], mid-circuit measurement[54, 55, 52, 53, 56, 7], and alternate addressing architectures[20, 57, 58, 7]. The advanced optical control techniques we demonstrate here also extend directly to alternative atomic species[20, 37, 59, 7], as well as multi-qubit ($N>2$) gates[60], which can further optimize logical operations[61].